g of Al (s) at 98° C is placed in 50 g of H2O(0), the final temperature i was the initial temperature of the water ?
Answer #1

-Qhot = Qcold

-Qal = Qwater

-m*C*(Tf-Tal) = mw*Cpw*(Tf-Twater)

substitute

-20*0.90*(26.5-98) = 50*4.184*(26.5-Twater)

-20*0.90*(26.5-98) /( 50*4.184) - 26.5 = -Tw

tw = -20.34

Twate rinitially = 20.34°C